OpenStudy (anonymous):

Could you give me an example of the difference between ODEs and PDEs? We got stuck with a terrible Diff Eqns book that uses no examples.

OpenStudy (anonymous):

This is ODE, only one variable y(x)=y′(x) solution y(x)= Ce^x y(x,y)=yx′(x,y) solution y(x,y)= f(y )e^x

OpenStudy (anonymous):

This is ODE, only one variable y(x)=y′(x) solution y(x)= Ce^x This is PDE;more than one variable \[y(x,y)=y_x′(x,y)\] solution y(x,y)= f(y )e^x

OpenStudy (asnaseer):

@imranmeah91 is correct. You may interested in these online references on these topics: 1) http://homepage.mac.com/nshoffner/nsh/ODE/Def_Class.html 2) http://tutorial.math.lamar.edu/Classes/DE/Definitions.aspx

OpenStudy (anonymous):

@imranmeah91, so the ODE would be written out as dy/dx, while the PDE might be written as y(d/dx)?